OpenStudy (anonymous):

what is 13/20 written as a decimal?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=13%2F20

OpenStudy (anonymous):

multiply both the top and bottom by 5 you would get \[\frac{13}{20}*\frac{5}{5}\]\[=\frac{65}{100}\]

OpenStudy (anonymous):

0.65

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

thx ill give u a medal

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

here u go

OpenStudy (anonymous):

thank you!